# Hadamard Attention Recurrent Transformer: A Strong Baseline for Stereo Matching Transformer

### v2

@@ Abstract
In light of the advancements in transformer technology, extant research posits the construction of stereo transformers as a potential solution to the binocular stereo matching challenge. However, constrained by the low-rank bottleneck and quadratic complexity of attention mechanisms, stereo transformers still fail to demonstrate sufficient nonlinear expressiveness within a reasonable inference time. The lack of focus on key homonymous points renders the representations of such methods vulnerable to challenging conditions, including reflections and weak textures. Furthermore, a slow computing speed is not conducive to the application. To overcome these difficulties, we present the Hadamard Attention Recurrent Stereo Transformer (HART) that incorporates the following components: 1) For faster inference, we present a Hadamard product paradigm for the attention mechanism, achieving linear computational complexity. 2) We designed a Dense Attention Kernel (DAK) to amplify the differences between relevant and irrelevant feature responses. This allows HART to focus on important details. DAK also converts zero elements to non-zero elements to mitigate the reduced expressiveness caused by the low-rank bottleneck. 3) To compensate for the spatial and channel interaction missing in the Hadamard product, we propose MKOI to capture both global and local information through the interleaving of large and small kernel convolutions. Experimental results demonstrate the effectiveness of our HART. In reflective area, HART ranked 1st on the KITTI 2012 benchmark among all published methods at the time of submission. Code is available at https://github.com/ZYangChen/HART.
Index Terms:
Hadamard Attention, Efficient Transformer, Stereo Matching.
publicationid: pubid: $0$–$00$ 2025 IEEE
I Introduction
Stereo matching represents a fundamental challenge in the field of computer vision, with a range of practical applications including autonomous driving, remote sensing, robotics and augmented reality [1, 2, 3, 4]. The goal of stereo matching is to compute the pixel-by-pixel correspondences between two images, and to generate a disparity map that can be translated into depth using the settings of the stereo camera system [5]. In practical, it requires both high accuracy and speed. Consequently, the attainment of efficient stereo matching represents a crucial objective.
@@ I Introduction
Since the development of deep learning, numerous stereo matching Convolutional Neural Networks (CNNs) have been presented. Recently, recurrent-based methods have shown promising performance on stereo matching benchmarks. RAFT-Stereo [6] uses a Recurrent All-pairs Field Transform (RAFT) [7] paradigm based on multi-level convolutional Gated Recurrent Units (convGRU). IGEV-Stereo [8] proposes the Combined Geometry Encoding Volume (CGEV) instead of the All-pairs Correlation (APC) Volume [6] in RAFT-Stereo. MoCha-Stereo [9] mining the recurrent feature in the feature channels and achieves state-of-the-art (SOTA) performance. However, the limited receptive field restricts the learning of long-range information. Recurrent-based methods face these challenges in matching local areas with unclear semantics, such as reflective regions and thin object areas.
The development of Vision Transformer (ViT) [10] has addressed the issue of acquiring long-range dependencies. STTR [11] uses a transformer architecture, replacing cost volume construction with dense pixel matching from a sequence-to-sequence correspondence perspective. ELFNet [12] combines stereo transformers with a multi-scale cost volume. However, existing stereo transformers still have certain limitations. As show in Fig. 1, stereo transformers are still constrained by the quadratic-level complexity of attention, limiting the speed of inference. They are challenging to meet the efficient computational requirements of scenarios such as autonomous driving. Moreover, vanilla self-attention (SA) also falls into a low-rank bottleneck [13]. This implies that stereo transformers have limited nonlinear expressive power, making them prone to being affected by ill-posed regions, such as reflections and weak textures, during the localization of homonymy points.
@@
1) To achieve an efficient design for stereo transformers, we propose a strategy of computing attention using the Hadamard product. This enables attention computation to be controlled within linear time complexity.
2) To circumvent the low-rank bottleneck inherent in vanilla SA, we design a Dense Attention Kernel (DAK) function, instead of using the SoftMax activation. DAK allows the rank of attention to have the potential to increase further.
3) To realize spatial and channel interactions, Multi Kernel & Order Interaction (MKOI) module is presented to learn global and local features.
4) HART ranked 1st on the KITTI 2012 reflective leaderboard and TOP 10 among all published methods at the time of submission. Furthermore, the attention structure of HART is scalable. HART has the potential to become a new baseline for stereo transformers.
II Related Work
II-A Recurrent-based Stereo CNNs
@@
Despite the leading performance of such methods, recurrent-based schemes typically use convolutional operations with relatively small receptive fields. They are not conducive to learning global information [10]. This makes recurrent-based CNNs challenging in areas where the semantic interpretation cannot be determined from local geometric information alone.
II-B Linear Transformers
@@
II-C Stereo Transformers
Benefiting from the global learning capability of ViT [10], stereo transformers [11, 12, 21] deliver excellent performance. STTR [11] constructs a transformer module with alternating self-attention and cross-attention to compute feature descriptors. CroCo-Stereo [21] stacks vanilla SA and MLP-designed encoder and decoder.
Though the design of the transformer allows for long-range matching, it is a computationally expensive model. SA has quadratic time complexity. As a result, existing stereo transformers generally require longer inference times. Currently, there are efforts to reduce the time complexity of stereo transformers. DLNR [22] borrows ideas from the attention in Restormer [23] and the recurrent update in RAFT-Stereo [6]. It deviates from the vanilla spatial feature matmul product and instead designs a channel feature matmul product. However, like other recurrent-based methods, DLNR lacks the spatial interactions of vanilla SA. In addition, although the number of channels $c$, is much smaller than $h\times w$, $c^{2}$ is still much larger than $h\times w$ in DLNR. This means that the SA computation of DLNR still has a time complexity of $O(c^{2})$.
In summary, stereo transformers still face two main challenges: 1) High computational cost of SA; 2) Inaccurate detail due to low-rank bottleneck.
@@ II-C Stereo Transformers
III Method
III-A Overview
We designed HART, a strong stereo transformer baseline via Linear time Hadamard product. This approach speeds up inference. Moreover, we believe that improving the matching performance of Stereo Transformers hinges on: 1) overcoming the low-rank bottleneck of SA; 2) ensuring interaction between spatial and channel features.
To further meet the accuracy requirements mentioned above, HART incorporates a series of new components made by Hadamard product: 1) DAK is designed to learn attention matrix with all non-zero elements, ensuring that HART has the potential for continued rank increase. Additionally, DAK emphasizes important features by assigning them larger response values. 2) MKOI is proposed to supplement the spatial and cross-channel interaction capabilities lacking in the Hadamard product.
@@ III-A Overview
The details of the above designs are described in this section. To complete the HART workflow, we introduce techniques for iterative disparity updating and correlation computation as part of the decoder. The full architecture of HART is portrayed in Fig. 2. The Hadamard Product Self-Attention (HPSA) process is described in Sections III-B and III-C. The Simple Gate Feed Forward (SGFF) mechanism is detailed in Section III-D. HPSA and SGFF form the encoder of HART, while the decoder of HART is outlined in Section III-E. Inspired by STTR [11] and IGEV-Stereo [8], we supervised HART by the Equ. 1, where $\gamma=0.9$, $N$ denotes the total number of iterations, $d_{0}$ represents the initial disparity computed from correlation volume in our IGEV Decoder, and the loss for the initial disparity is calculated using smooth L1. L1 loss is used for subsequent iteration processes.
$L=Smooth_{L1}(d_{0}-d_{gt})+\sum_{i=1}^{N}\gamma^{N-i}||d_{i}-d_{gt}||_{1}$ (1)
III-B Caculation of Attention Matrix
@@
Although stereo transformers can achieve good results in long-range matching, this architecture is still not efficient enough. First, existing stereo transformers lack the non-linearity to convey complex scene detail. This is due to the low-rank bottleneck. Secondly, the quadratic complexity of the attention mechanism makes it difficult for transformers to achieve efficient performance. To achieve linear time complexity, HPSA is designed to be used here. We also implemented a series of designs to prevent HPSA from falling into the low-rank bottleneck.
III-B1 Hadamard Product between Query and Key
@@
Vanilla SA [10] in the Stereo Transformer [11, 12, 21] relies on matmul product to compute the attention matrix $A$. However, this computational approach exhibits a quadratic level of complexity. $n=h\times w$ here, $h$ denotes the height, $w$ means width of the feature map. $Q,K,V,A$ stand for query, key, value and attention matrix respectively. In existing Stereo Transformers, it holds that $n>>c$. Consequently, the vanilla SA has a time complexity of $O(n^{2})$. Despite the introduction of channel-wise SA [23, 22] for Stereo Transformer, the condition $c^{2}>>n$ still prevails. To make attention more efficient, we have developed the HPSA as illustrated in Equ. 2.
$A=||Q||_{2}\odot||K||_{2}$ (2)
@@
III-B2 Dense Attention Kernel (DAK)
Attention matrix $A$ comprises only $n\times c$ parameters, and the $Q$ and $K$ matrices together consist of $2nc$ parameters. This implies that, in theory, each parameter can be adequately represented. To achieve this, we aim for all parameters to maintain a non-zero state. Following this rationale, we have designed the Dense Attention Kernel (DAK), as depicted in Equ. 4. This non-zero state can alleviate the low-rank bottleneck [13], ensuring that pixel features can be expressed using a more diverse set of parameters.
@@ III-B2 Dense Attention Kernel (DAK)
Different from ELU activation [24], our kernel function differs significantly in its underlying principles: 1) Our designed DAK aims to be utilized as an kernel function for attention. 2) In comparison to the SoftMax utilized in vanilla SA and ELU used as activate function, our DAK always satisfied $DAK(A)\textgreater 0$. DAK suppresses the occurrence of silent neurons. Because zero elements in the matrix may become non-zero, there exists the theoretical possibility of an increase in the rank of the matrix. This implies that attention can have stronger non-linear expression capabilities. This design choice also involves assigning a smaller activation value to relatively less important dimensions. This serves as a weight in the description of the characteristics for the Value $V$.
We can ensure that each dimension is positively correlated with the weights here. Because the feature values of $A$ at a given pixel are described by the channel features of that pixel. 3) In addition to amplifying the differences in attentional response scores, DAK shows more robust expressiveness than ELU. This is attributed to the fact that DAK can be decoupled into a process involving Equ. 5.
@@
Furthermore, in contrast to the SoftMax function, which confines values to the $[0,1]$ interval, DAK extends the allocation of attention weights to the higher-dimensional interval $(0,+\infty)$. The derivative properties of DAK determine its ability to further widen the attentional gap, thereby redirecting the attention to more important features. The broader range of attentional weighting contributes to a nuanced emphasis on key features. These ensure more diverse and focused feature representation, theoretically allowing for more precise matching.
III-C Multi Kernel & Order Interaction (MKOI)
Vanilla self-attention [10] satisfied $SA(Q,K,V)=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$. This process involves spatial and channel interactions. Vanilla Hadamard product cannot achieve this. To improve the spatial and channel interaction of HPSA, we propose the Multi Kernel & Order Interaction (MKOI) module as a complement to HPSA.
As shown in Fig. 3, MKOI is a Hadamard product paradigm concerning the Attention matrix $A$ and the Value matrix $V$. To supplement the channel interaction capability, MKOI decouples features into four sets with channel numbers $[c,c/2,c/4,c/4]$, where $c$ represents the original number of channels. The rough semantic distributions of the image are already implied by lower order features [26]. Therefore, we use large kernel convolutions to extract global semantic distributions. Compared to applying large kernel convolutions to all channels [27, 28, 29], this extraction mode is obviously cheaper. Higher-order features can focus on local features, so we use 1$\times$1 and 3$\times$3 convolutions to extract details. Through this paradigm, MKOI can capture both long-range and short-range information.
@@ III-C Multi Kernel & Order Interaction (MKOI)
$HPSA(Q,K,V)=w^{2c\rightarrow c}_{1\times 1}([\underline{DAK(||Q^{\frac{c}{2^{m}}}||_{2}\odot||K^{\frac{c}{2^{m}}}||_{2})}$
$\underline{\odot w^{c\rightarrow\frac{c}{2^{m}}}_{s\times s}(V),DAK(||Q^{\frac{c}{4}}||_{2}\odot||K^{\frac{c}{4}}||_{2})\odot w^{c\rightarrow\frac{c}{4}}_{7\times 7}(V)}]_{m=\{0,1,2\}})$ (7)
@@
Overall, the HPSA combined with MKOI can be expressed by Equ. 7 and Fig. 1 (b). MKOI uses 1 $\times$ 1 convolutions to compress and interact channels, producing $V_{i+1}$ as input to the next transformer block. $[\underline{\dots}]$ means concat along the feature channel dimension here, $s$ denotes the kernel size and $s=2m+1$.
@@
III-D Simple Gate Feed Forward (SGFF)

### v4

@@ Abstract
Constrained by the low-rank bottleneck inherent in attention mechanisms, current stereo matching transformers suffer from limited nonlinear expressivity, which renders their feature representations sensitive to challenging conditions such as reflections. To overcome this difficulty, we present the Hadamard Attention Recurrent Stereo Transformer (HART). HART includes a novel attention mechanism that incorporates the following components: 1) The Dense Attention Kernel (DAK) maps the attention weight distribution into a high-dimensional space over (0, $\infty$). By removing the upper bound constraint on attention weights, DAK enables more flexible modeling of complex feature interactions. This reduces feature collinearity. 2) The Multi Kernel & Order Interaction (MKOI) module extends the attention mechanism by unifying semantic and spatial knowledge learning. This integration improves the ability of HART to learn features in binocular images. Experimental results demonstrate the effectiveness of our HART. In reflective area, HART ranked 1st on the KITTI 2012 benchmark among all published methods at the time of submission. Code is available at here.
keywords:
[…]
highlights
Hadamard Attention Recurrent Stereo Transformer (HART) is an efficient stereo transformer based on the hadamard product. The attention mechanism of HART is designed to be scalable, positioning HART as a potential new baseline for stereo transformer models.
To enhance matching performance under challenging conditions, we propose the Dense Attention Kernel (DAK) to address the low-rank bottleneck in transformers. DAK improves the nonlinear expressiveness of the model, enabling a better understanding of disparity relationships in complex scenarios.
To enhance spatial and channel interactions, we introduce the Multi-Kernel & Order Interaction (MKOI) module, designed to capture both global and local features.
HART ranked 1st on the KITTI 2012 reflective leaderboard and and secured a position in the Top ten on the Middlebury benchmark among all published methods at the time of submission.
1 Introduction
Stereo matching represents a fundamental challenge in the field of computer vision, with a range of practical applications including autonomous driving, remote sensing, robotics and augmented reality [1, 2, 3]. The goal of stereo matching is to compute the pixel-by-pixel correspondences between two images, and to generate a disparity map that can be translated into depth using the settings of the stereo camera system [4]. In practical, it needs to achieve excellent matching capabilities in a variety of scenarios.
@@ 1 Introduction
Since the development of deep learning, numerous stereo matching Convolutional Neural Networks (CNNs) have been presented. Recently, recurrent-based methods have shown promising performance on stereo matching benchmarks. RAFT-Stereo [5] uses a Recurrent All-pairs Field Transform (RAFT) [6] paradigm based on multi-level convolutional Gated Recurrent Units (convGRU). IGEV-Stereo [7] proposes the Combined Geometry Encoding Volume (CGEV) instead of the All-pairs Correlation (APC) Volume [5] in RAFT-Stereo. MoCha-Stereo [8] mining the recurrent feature in the feature channels and achieves state-of-the-art (SOTA) performance. While the limited receptive field restricts the learning of long-range information.
The development of Vision Transformer (ViT) [9] has addressed the issue of acquiring long-range dependencies. STTR [10] uses a transformer architecture, replacing cost volume construction with dense pixel matching from a sequence-to-sequence correspondence perspective. ELFNet [11] combines stereo transformers with a multi-scale cost volume. However, existing stereo transformers still have certain limitations.
As show in Fig. 1, stereo transformers are still constrained by the quadratic-level complexity of attention, limiting the speed of inference. They are challenging to meet the efficient computational requirements of scenarios such as autonomous driving. Moreover, vanilla self-attention (SA) also falls into a low-rank bottleneck [12]. Specifically, each query matrix $Q$ and key matrix $K$ for every attention head jointly comprise only $\frac{2nc}{d}$ parameters, where $n=h\times w$ represents the spatial dimension, $h$ and $w$ denote the height and width of feature maps separately, $c$ represents the channel dimension, $d$ means the number of attention heads. Meanwhile, the attention matrix $A=Q\times K$, where the number of parameters is $n^{2}$, satisfies $\frac{2nc}{d}<<n^{2}$. This implies that stereo transformers have limited nonlinear expressive power, making them prone to being affected by ill-posed regions, such as reflections and weak textures.
@@
1) To achieve an efficient architecture, we propose a novel stereo transformer pipeline based on the Hadamard product. This approach ensures that the computation is performed with linear time complexity.
2) To enhance matching performance under challenging conditions, we propose the Dense Attention Kernel (DAK) to address the low-rank bottleneck in transformers. DAK improves the nonlinear expressiveness of the model, enabling a better understanding of disparity relationships in complex scenarios.
3) To enhance spatial and channel interactions, we introduce the Multi-Kernel & Order Interaction (MKOI) module, designed to effectively capture both global and local features.
4) HART ranked 1st on the KITTI 2012 reflective leaderboard and and secured a position in the Top ten on the Middlebury benchmark among all published methods at the time of submission. Moreover, the attention mechanism of HART is designed to be scalable, positioning HART as a potential new baseline for stereo transformer models.
2 Related Work
2.1 Recurrent-based Stereo CNNs
@@
Despite the leading performance of such methods, recurrent-based schemes typically use convolutional operations with relatively small receptive fields. They are not conducive to learning global information [9, 16]. This makes recurrent-based CNNs challenging in areas where the semantic interpretation cannot be determined from local geometric information alone.
2.2 Stereo Transformers
Benefiting from the global learning capability of ViT [9], stereo transformers [10, 11, 17] deliver excellent performance. STTR [10] constructs a transformer module with alternating self-attention and cross-attention to compute feature descriptors. ELFNet [11] utilizes self-attention transformer to learn global features as part of estimating evidence-based disparities. CroCo-Stereo [17] stacks SA and MLP-designed encoder and decoder. Since both approaches incorporate vanilla self-attention (SA) within their transformer blocks, existing stereo transformers struggle to match ill-posed regions such as reflective surfaces and thin structures.
Though the design of the transformer allows for long-range matching [18, 9, 19], it is a computationally expensive model. SA has quadratic time complexity. As a result, existing stereo transformers generally require longer inference times. Currently, there are efforts to reduce the time complexity of stereo transformers. DLNR [20] and RetinaStereo [21] borrow ideas from the attention in Restormer [22, 23] and the recurrent update in RAFT-Stereo [5]. It deviates from the vanilla spatial feature matmul product and instead designs a channel feature matmul product. However, like other recurrent-based methods, DLNR and RetinaStereo lack the spatial interactions of vanilla SA. In addition, although the number of channels $c$, is much smaller than $h\times w$, $c^{2}$ is still much larger than $h\times w$ in DLNR and RetinaStereo. This means that the SA computation of DLNR and RetinaStereo [20, 21] still have a time complexity of $O(c^{2})$.
In summary, stereo transformers still face two main challenges: 1) Inaccurate detail due to low-rank bottleneck; 2) High computational cost of SA. We hypothesize that the low-rank bottleneck inherent in vanilla SA restricts the expressive capacity of stereo transformers. The limited nonlinear expressiveness of SA makes it difficult for the model to capture homonymy point correspondences in ill-posed regions. Additionally, to improve computational efficiency, we aim to reduce the complexity of the attention mechanism from $O(n^{2})$ to $O(n)$.
2.3 Efficient Transformers for Vision Tasks
@@
3 Method
3.1 Overview
We designed Hadamard Attention Recurrent Transformer (HART), a strong baseline for stereo transformers. We believe that improving the matching performance of stereo transformers hinges on overcoming the low-rank bottleneck [12] of SA, and incorporates a series of new components: 1) DAK broadens the range of activation values, thereby reducing feature similarity. This mechanism alleviates the expressiveness degradation caused by the low-rank bottleneck, which facilitates the HART in matching homonymous points within ill-posed regions. 2) MKOI is proposed to supplement the spatial and cross-channel interaction capabilities lacking in the Hadamard product.
@@ 3.1 Overview
DAK and MKOI jointly constitute the novel attention mechanism we propose, referred to as Hadamard Product Self-Attention (HPSA). The HPSA process is described in Sections 3.2 and 3.3. To complete the HART workflow, we introduce techniques for iterative disparity updating and correlation computation as part of the decoder. The full architecture of HART is portrayed in Fig. 2. The Simple Gate Feed Forward (SGFF) mechanism is detailed in Section 3.4. HPSA and SGFF form the encoder of HART, the decoder of HART is outlined in Section 3.5. Inspired by STTR [10] and IGEV-Stereo [7], we supervised HART by the Equ. 1, where $\gamma=0.9$, $N$ denotes the total number of iterations, $d_{0}$ represents the initial disparity computed from correlation volume in our IGEV Decoder, and the loss for the initial disparity is calculated using smooth L1. L1 loss is used for subsequent iteration processes.
[…]
3.2 Caculation of Attention Matrix
@@
Although stereo transformers can achieve good results in long-range matching, this architecture is still not efficient enough. First, stereo transformers often produce incorrect matches for homonymous points in ill-posed regions. We attribute this issue to the low-rank bottleneck inherent in vanilla SA. Secondly, the quadratic complexity of the attention mechanism makes it difficult for transformers to achieve efficient performance. To achieve linear time complexity, HPSA is designed to be used here. We implemented a series of designs to prevent HPSA from falling into the low-rank bottleneck.
3.2.1 Hadamard Product between Query and Key
@@
Vanilla SA [9] in the stereo transformer [10, 11, 17] relies on matmul product to compute the attention matrix $A$. However, this computational approach has a time complexity of $O(n^{2})$. $n=h\times w$ here, $h$ denotes the height, $w$ means width of the feature map. $Q,K,V,A$ stand for query, key, value and attention matrix respectively. In existing stereo transformers, it holds that $n>>c$, where $c$ means the number of feature channels. The above conclusion is a proof that the vanilla SA exhibits a quadratic level of complexity. Despite the introduction of channel-wise SA [22, 20] for stereo transformers, the condition $c^{2}>>n$ still prevails. To improve the computational efficiency of attention, we compute the attention matrix using the Hadamard product, as illustrated in Equ. 2.
[…]
@@
3.2.2 Dense Attention Kernel (DAK)
Attention matrix $A$ comprises only $n\cdot c$ parameters, and the $Q$ and $K$ matrices together consist of $2\cdot n\cdot c$ parameters. This implies that, in theory, each parameter can be adequately represented. To achieve this, we aim for all parameters to maintain a non-zero state. Following this rationale, we have designed the Dense Attention Kernel (DAK), as depicted in Equ. 4.
@@ 3.2.2 Dense Attention Kernel (DAK)
[…]
In addition to amplifying the differences in attentional response scores, DAK also shows more robust expressiveness than ELU. This is attributed to the fact that DAK can be decoupled into a process involving Equ. 5.
@@
These ensure more diverse feature representation, theoretically allowing for more precise matching.
3.3 Multi Kernel & Order Interaction (MKOI)
Vanilla self-attention (SA) [9] satisfied $SA(Q,K,V)=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$. This process involves spatial and channel interactions. Vanilla Hadamard product cannot achieve this. To improve the spatial and channel interaction of HPSA, we propose the Multi Kernel & Order Interaction (MKOI) module as a complement to HPSA.
As shown in Fig. 3, MKOI is a Hadamard product paradigm concerning the Attention matrix $A$ and the Value matrix $V$. To supplement the channel interaction capability, MKOI decouples features into four sets with channel numbers $[c,c/2,c/4,c/4]$, where $c$ represents the original number of channels. The rough semantic distributions of the image are already implied by lower order features [29]. Therefore, we use large kernel convolutions to extract global semantic distributions. Compared to applying large kernel convolutions to all channels [30], this extraction mode is obviously cheaper. Higher-order features can focus on local features, so we use $\times1$ and $\times3$ convolutions to extract details. Through this paradigm, MKOI can capture both long-range and short-range information.
@@ 3.3 Multi Kernel & Order Interaction (MKOI)
[…]
$\underline{\odot w^{c\rightarrow\frac{c}{2^{m}}}_{s\times s}(V)}]_{m=\{0,1,2\}})$ (7)
@@
Overall, the HPSA combined with MKOI can be expressed by Equ. 7 and Fig. 1 (b). MKOI uses 1 $\times$ 1 convolutions to compress and interact channels, producing $V_{i+1}$ as input to the next transformer block. $\underline{\dots}$] means concat along the feature channel dimension here, $s$ denotes the kernel size and $s=2m+3$.
@@
3.4 Simple Gate Feed Forward (SGFF)